# Learning Motion Blur Robust Vision Transformers with Dynamic Early Exit for Real-Time UAV Tracking

### v1

@@ IV-G3 Study on weighting the loss for learning motion blur robust ViTs (MBRV)
To observe how the weight $\rho$ of the loss $\mathcal{L}_{br}$ influence the performance, we train BDTrack-DeiT with various $\rho$ values ranging from $0.5\times 10^{-4}$ to $5.0\times 10^{-4}$. The evaluation results assessed across all five datasets are shown in Table V. Obviously, BDTrack-DeiT achieves averagely the optimal performance at $\rho$ = $1.0\times 10^{-4}$, with 84.2% Avg. Prec. and 64.4% Avg. Succ.. Additionally, we can observe that the second and third-best performances across these datasets are scattered both above and below the value of $1.0\times 10^{-4}$. These fluctuations may be attributed to the inherent differences between these datasets. When $\rho$ is set as $4.0\times 10^{-4}$, BDTrack-DeiT achieves the lowest average performance, with 82.6% Avg. Prec. and 63.2% Avg. Succ. This indicates that selecting an appropriate weight $\rho$ is crucial for achieving optimal tracking performance. An ill-suited $\rho$ can adversely affect the training process, leading to suboptimal results. The performance variations observed with different $\rho$ values highlight the sensitivity of the tracking algorithm to this parameter. Therefore, careful tuning of $\rho$ is necessary to ensure that the model performs effectively across diverse datasets and scenarios.
IV-G4 Study on weighting the proposed block sparsity loss
In order to provide a detailed analysis of the impact of the proposed block sparsity loss $\mathcal{L}_{spar}$ on the performance, Table VI presents the evaluation results of BDTrack-DeiT trained with various block sparsity loss weight $\gamma$ values ranging from $0.5\times 10^{3}$ to $1.5\times 10^{3}$. Note that during training, the weight $\rho$ of the loss $\mathcal{L}_{br}$ is set to $1.0\times 10^{-4}$. As shown in the table, BDTrack-DeiT achieves the optimal performance when the $\gamma$ is set to $1.0\times 10^{3}$. When $\gamma$ is increased to $1.5\times 10^{3}$, BDTrack-DeiT shows a significant drop, achieving the lowest average performance with a 1.5% decrease in Avg. Prec. and a 1.4% decrease in Avg. Succ. The experimental results demonstrate that the choice of weight $\gamma$ significantly impacts tracking performance, and improper selection of this parameter can lead to suboptimal performance.
IV-G5 Application to SOTA trackers
@@
To demonstrate the general applicability of our approach, we integrate the proposed MBRV and DEEM into three SOTA trackers: ARTrack [80], GRM [69], and DropTrack [81]. Note that we substitute their original backbones with the tiny ViT, i.e., ViT-Tiny [65], to save training time. The evaluation results on five benchmarks are presented in Table VII. As mentioned previously, we provide only the speed of the baseline to eliminate potential nuances arising from randomness. We can observe that the inclusion of MBRV leads to noticeable improvements in both Prec. and Succ. for all baselines. Specifically, ARTrack, GRM, and DropTrack experience increases of 1.8%, 1.3%, and 1.7% in Avg. Prec., respectively, while their Avg. Succ. grows by 1.3%, 1.0%, and 1.1%, respectively. When DEEM is further integrated, we consistently observe improvements in GPU speeds, along with only slight decreases in Prec. and Succ. More specifically, the GPU speeds of all baselines demonstrate increases of more than 14.0%, with tracking performance dropping by less than 0.6%. These results confirm the applicability of our method in enhancing the efficiency of existing ViT-based trackers without significantly compromising tracking performance, justifying its generality.
TABLE VII: Application of our MBRV and DEEM to SOTA trackers. Note that precision and success rate are shown in the form of (Prec., Succ.)
Tracker MBRV DEEM DTB70 UAVDT VisDrone2018 UAV123 UAV123@10fps Avg. Avg.FPS
ARTrack[80] (79.6, 61.2) (74.7, 52.8) (78.6, 60.1) (80.4, 63.1) (80.1, 63.3) (78.6, 60.1) 78.3
$\checkmark$ (80.1, 62.0) (76.7, 54.7) (79.8, 60.6) (81.6, 63.9) (83.9, 66.0) (80.4↑1.8, 61.4↑1.3) -
$\checkmark$ $\checkmark$ (79.2, 61.5) (76.0, 54.1) (79.4, 60.3) (80.9, 63.6) (83.4, 65.7) (79.8↑1.2, 61.0↑0.9) 89.3↑14%
GRM [69] (81.8, 63.6) (78.1, 56.9) (85.5, 64.8) (84.6, 66.5) (82.6, 65.3) (82.5, 63.4) 224.3
$\checkmark$ (83.1, 65.2) (80.4, 58.4) (86.2, 65.3) (85.4, 67.1) (84.0, 66.0) (83.8↑1.3, 64.4↑1.0) -
$\checkmark$ $\checkmark$ (82.9, 64.9) (79.9, 58.1) (85.5, 64.8) (85.0, 66.8) (83.5, 65.7) (83.4↑0.9, 64.1↑0.7) 261.1↑16%
DropTrack [81] (83.4, 64.7) (79.4, 57.5) (81.9, 61.8) (84.8, 66.2) (82.5, 65.1) (82.4, 63.1) 214.6
$\checkmark$ (84.2, 65.3) (81.6, 58.9) (85.1, 64.2) (85.4, 66.6) (84.0, 66.1) (84.1↑1.7, 64.2↑1.1) -
$\checkmark$ $\checkmark$ (83.5, 64.9) (81.0, 58.4) (84.7, 64.3) (85.0, 66.3) (83.7, 65.8) (83.6↑1.2, 63.9↑0.8) 253.8↑18%
@@
IV-G6 Qualitative Results
Our BDTrack enhances the robustness of Vision Transformers (ViTs) against motion blur by enforcing invariance in the feature representation of the target. To intuitively demonstrate the effectiveness of the proposed components, Fig. 7 and Fig. 8 visualize several example attention maps of the search images and example feature maps of the template images generated by BDTrack with and without the proposed components (i.e., MBRV and DEEM). Note that in these examples the targets are subjected to motion blur. For convenience, a ’*’ suffix is added to BDTrack-DeiT to denote the baseline that excludes any of the proposed components. A ’+’ suffix indicates the inclusion of the MBRV component in the baseline, while a ’++’ suffix signifies the incorporation of both the MBRV and DEEM components into the baseline.
Example attention maps of the search images. In Fig. 7, the first row shows the original search images from two example sequences (i.e., SpeedCar4 and uav1), one from DTB70 [64] and the other from UAV123 [61]. The images in the second, third, and fourth rows are the corresponding attention maps generated by BDTrack-DeiT*, BDTrack-DeiT+, and BDTrack-DeiT++ (i.e., BDTrack-DeiT), respectively. As can be seen, incorporating the MBRV component (i.e., BDTrack-DeiT+) and both the MBRV and DEEM components (i.e., BDTrack-DeiT++) yields very similar results. Specifically, their visual focus on the targets is more pronounced compared to the baseline tracker (i.e., BDTrack-DeiT*). This demonstrates that the MBRV component is able to enhance the tracker’s ability to concentrate on the target, especially under motion blur conditions, thereby improving overall tracking performance.
@@ IV-G6 Qualitative Results
IV-H Real-World Test
@@
V Conclusions
In this work, our focus was on exploring efficient and motion blur robust Vision Transformers (ViTs) within a unified framework for real-time UAV tracking. We achieved this by tailoring the framework into an adaptive computation paradigm that dynamically early exits Transformer blocks. Moreover, in addressing the prevalent challenges associated with motion blur in UAV tracking, our approach focuses on learning motion blur robust Vision Transformers (ViTs). This is achieved by enforcing invariance in the feature representation of the target against simulated motion blur. Remarkably, our method stands out for its simplicity, making it easily integrable or adaptable into other tracking frameworks without necessitating additional architectures. Extensive experiments across five challenging UAV benchmarks were conducted to validate the effectiveness of our proposed method. The experimental results demonstrate that our BDTrack-DeiT achieves state-of-the-art performance in real-time UAV tracking.
Acknowledgments
Thanks to the support by the Guangxi Natural Science Foundation (Grant No. 2024GXNSFAA010484), the Guangxi Science and Technology Base and Talent Special Project (No. GKAD22035127), Guangxi Key Technologies R&D Program (Grant No. AB23049001), the National Natural Science Foundation of China (Nos. 62206123, 62066042, 62262011, 62176170, and 61971005), and the Sichuan Province Key Research and Development Project (No. 2020YJ0282), this work has been made possible.

### v2

@@ 4.7.2 Study on weighting the loss for learning motion blur robust ViTs (MBRV)
To observe how the weight $\rho$ of the loss $\mathcal{L}_{br}$ influence the performance, we train BDTrack-DeiT with various $\rho$ values ranging from $0.5\times 10^{-4}$ to $5.0\times 10^{-4}$. The evaluation results assessed across all four datasets are shown in Table 5. Obviously, BDTrack-DeiT achieves averagely the optimal performance at $\rho$ = $1.0\times 10^{-4}$, with 84.2% Avg. Prec. and 64.4% Avg. Succ.. Additionally, we can observe that the second and third-best performances across these datasets are scattered both above and below the value of $1.0\times 10^{-4}$. These fluctuations may be attributed to the inherent differences between these datasets. When $\rho$ is set as $0.8\times 10^{-4}$, BDTrack-DeiT achieves the lowest average performance, with 82.5% Avg. Prec. and 63.1% Avg. Succ. This indicates that selecting an appropriate weight $\rho$ is crucial for achieving optimal tracking performance. An inappropriate $\rho$ can adversely affect the training process, leading to suboptimal results. The performance variations observed with different $\rho$ values highlight the sensitivity of the tracking algorithm to this parameter. Therefore, careful tuning of $\rho$ is necessary to ensure that the model performs effectively across diverse datasets and scenarios.
4.7.3 Application to SOTA trackers
@@
To demonstrate the general applicability of our approach, we integrate the proposed MBRV and DEEM into two SOTA trackers: GRM [89] and DropTrack [90]. Note that we substitute their original backbones with the tiny ViT, i.e., ViT-Tiny [74], to save training time. The evaluation results on three benchmarks are presented in Table 6. As mentioned previously, we provide only the speed of the baseline to eliminate potential nuances arising from randomness. We can observe that the inclusion of MBRV leads to noticeable improvements in both Prec. and Succ. for all baselines. Specifically, GRM and DropTrack experience increases of 1.4% and 2.3% in Avg. Prec., respectively, while their Avg. Succ. grows by 0.9% and 1.6%, respectively. When DEEM is further integrated, we consistently observe improvements in GPU speeds, along with only slight decreases in Prec. and Succ. More specifically, the GPU speeds of all baselines demonstrate increases of more than 17.0%, with tracking performance dropping by less than 0.5%. These results confirm the applicability of our method in enhancing the efficiency of existing ViT-based trackers without significantly compromising tracking performance, justifying its generality.
Table 6: Application of our MBRV and DEEM to two SOTA trackers, with their original backbones replaced by ViT-Tiny [74] to reduce training time.
Tracker MBVR DEEM UAVDT VisDrone2018 UAV123@10fps Avg Avg.FPS
Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ.
GRM [89] 78.1 56.9 85.5 64.8 82.6 65.3 82.1 62.3 218.6
$\checkmark$
80.4
58.4
86.2
65.3
84.0
66.0
83.5
63.2
$\checkmark$ $\checkmark$ 79.9 58.1 85.5 64.8 83.5 65.7 83.0 62.9
257.3
↑17.8%
DropTrack [90] 79.4 57.5 81.9 61.8 82.5 65.1 81.3 61.5 211.4
$\checkmark$
81.6
58.9
85.1
64.2
84.0
66.1
83.6
63.1
$\checkmark$ $\checkmark$ 81.0 58.4 84.7 63.8 83.7 65.8 83.1 62.7
248.6
↑17.5%
Table 7: Comparison of performance and speed between our DEEM and various early exit methods on three UAV tracking datasets. Note that the baseline represents the enhanced tracker, which already includes the proposed MBVR.
UAVDT VisDrone2018 UAV123@10fps Avg.
Tracker Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. Avg.FPS
baseline 84.5 61.4 86.1 65.1 84.2 66.2 84.9 64.2 228.7
+ DEEM 84.1 61.0 85.2 64.3 83.5 65.9 84.3 63.7 279.4↑22.2%
+ PABEE [34] 81.7 58.5 84.1 63.8 82.1 63.8 82.6 62.0 298.1 ↑30.7%
+ ViT-EE [91] 81.1 58.2 82.5 61.7 83.0 65.6 82.2 61.8 318.7 ↑39.4%
+ LGViT [6] 82.7 59.2 83.1 62.0 84.1 66.2 83.3 62.5 264.5↑15.7%
4.7.4 Comparison with Existing Early Exit Methods
We have conducted a comparative experiments between our method and existing early exit mechanisms, including PABEE [34], ViT-EE [91], and LGViT [6]. The evaluation results are shown in Table 7. As observed, while these methods employ more complex architectures, they introduce additional complexity, struggle to balance speed and accuracy, and may exit prematurely, resulting in suboptimal performance in challenging tracking scenarios. Specifically, although PABEE and ViT-EE achieve faster speeds than our DEEM, their performance is inferior, with an average drop of over 2.0% in precision and success rate. Additionally, while LGViT achieves relatively better performance than PABEE and ViT-EE, it comes at the cost of even slower speeds. In contrast, with the inclusion of our DEEM, the baseline’s performance drops by less than 0.6% in average precision and success rate, while achieving a speedup of over 22%. These comparisons highlight the effectiveness of the proposed DEEM in UAV tracking.
@@ 4.7.4 Comparison with Existing Early Exit Methods
4.7.5 Qualitative Results
Our BDTrack improves ViT robustness to motion blur by enforcing target feature invariance. To illustrate the effectiveness of our components, Fig.6 and Fig.7 show example attention and feature maps generated by BDTrack with and without the proposed modules (i.e., MBRV and DEEM), under motion blur conditions. For clarity, we use suffixes to indicate model variants: ‘*’ denotes the baseline without our components, ‘+’ adds MBRV to the baseline, and ‘++’ includes both MBRV and DEEM.
Example attention maps of the search images. In Fig. 6, the first row shows the original search images from an example sequence (i.e., uav1 from UAV123 [60]). The images in the second, third, and fourth rows are the corresponding attention maps generated by BDTrack-DeiT*, BDTrack-DeiT+, and BDTrack-DeiT++ (i.e., BDTrack-DeiT), respectively. As can be seen, incorporating the MBRV component (i.e., BDTrack-DeiT+) and both the MBRV and DEEM components (i.e., BDTrack-DeiT++) yields very similar results. Specifically, their visual focus on the targets is more pronounced compared to the baseline tracker (i.e., BDTrack-DeiT*). This demonstrates that the MBRV component is able to enhance the tracker’s ability to concentrate on the target, especially under motion blur conditions, thereby improving overall tracking performance.
@@ 4.7.5 Qualitative Results
4.8 Real-World Test
@@
Figure 9: BDTrack limitations in complex tracking situations: Tracking errors occur when the object is fully occluded, indistinguishable from the surroundings, or occupies a very small region.
5 Limitations and future work
5.1 Limitations
We present several representative failure cases in Fig. 9. Although our BDTrack demonstrates strong performance in both efficiency and motion blur robustness, it still faces challenges in certain complex tracking scenarios. Specifically, the current design lacks an explicit mechanism for recovering target features after full occlusion, which may lead to identity loss or tracking failure once the target reappears (as shown in the top row). Moreover, the tracker’s performance degrades when dealing with extremely small or visually similar (as shown in the bottom row). This is primarily due to the limited spatial resolution of high-level features and the absence of fine-grained contrast enhancement, which affects the ability to distinguish the target from the background and increases the risk of drift.
5.2 Future work
To address these limitations, future work will focus on enhancing the tracker’s robustness in scenarios involving full occlusion and low visibility. One potential direction is to incorporate lightweight memory mechanisms that enable the tracker to re-identify the target after temporary disappearance. Additionally, we aim to improve the model’s sensitivity to fine-grained visual cues by integrating high-resolution refinement modules, thereby facilitating more accurate localization of small targets.
6 Conclusions
In this work, we presented a unified framework that explores both efficiency and motion blur robustness for real-time UAV tracking with Vision Transformers (ViTs). Our design introduces an adaptive computation paradigm via a dynamic early exiting mechanism, significantly reducing inference cost without compromising accuracy. Furthermore, we proposed a motion blur robust learning strategy that enforces feature invariance between clear and blurred templates, enhancing the tracker’s robustness under adverse conditions.
Extensive experiments conducted on four UAV tracking benchmarks demonstrate the effectiveness and versatility of our approach. In particular, BDTrack-DeiT achieves a more favorable speed-accuracy trade-off compared to existing state-of-the-art trackers. It outperforms lightweight models such as Aba-ViTrack [9], running approximately 1.6 times faster on GPU and 1.3 times faster on CPU, and surpasses deeper models like ROMTrack [77] and EVPTrack [78] with speed improvements of about 5 and 11 times, respectively, while achieving comparable or even better accuracy. The experimental results demonstrate that our BDTrack-DeiT achieves state-of-the-art performance in real-time UAV tracking.